### v1

@@ 5 Conclusion
In this work, we have presented an implementation of multi-jet merging in DIS with Pythia, using the Vincia antenna-shower framework. We have employed the CKKW-L and UMEPS algorithms with parton-level event input from Sherpa and cross-checked the results using hard processes from MadGraph5. We have studied the effects of different factorization and renormalization-scale choices along with different merging-scale prescriptions. We found that cross sections for $N$-jet production converge when we included events up to $N+1$ partons in the merging algorithms. Comparing the results to experimental data from the H1 collaboration, we find good agreement of differential multi-jet cross sections across a wide range of photon virtualities.
In case of CKKW-L merging we studied two different merging scales, a fixed one and a $Q^{2}$-dependent option referred to as dynamic merging scale. In case of the former the renormalization and factorization scales accounted also for the transverse momentum of the produced partons to partly capture the idea of the dynamic merging scale where the part of the phase space with high-$p_{\mathrm{T}}$ partons and low $Q^{2}$ would be mainly covered by the hard partons instead of parton shower. Indeed, we found that the resulting cross sections were very similar with these approaches. However, we found that the fixed merging scale lead to an increased sensitivity to the value of the merging scale compared to the dynamic setting. In case of UMEPS we could only use the fixed merging scale option. This algorithm lead to a more stable cross sections against the merging scale variations. However, the agreement with the H1 data was somewhat worse than with CKKW-L for inclusive and dijet production at high $p_{\mathrm{T}}$.
@@ Acknowledgements.
We would like to express our gratitude to Frank Krauss, Leif Lönnblad, and Olivier Mattelaer for their insightful discussions and valuable answers regarding this study. We would like to thank Stefan Höche for comments on the DIS merging implementation in Sherpa and help with the scale setting in the HPC-enabled version of Sherpa. We acknowledge grants of computer capacity from the Finnish Grid and Cloud Infrastructure (persistent identifier urn:nbn:fi:research-infras-2016072533). Part of the computations were carried out on the PLEIADES cluster at the University of Wuppertal, supported by the Deutsche Forschungsgemeinschaft (DFG, grant No. INST 218/78-1 FUGG) and the Bundesministerium für Bildung und Forschung (BMBF).

### v3

@@ 5 Conclusion
In this work, we have presented an implementation of multi-jet merging in DIS with Pythia, using the Vincia antenna-shower framework. We have employed the CKKW-L and UMEPS algorithms with parton-level event input from Sherpa and cross-checked the results using hard processes from MadGraph5. We have studied the effects of different factorization and renormalization-scale choices along with different merging-scale prescriptions. We found that cross sections for $N$-jet production started to level off when we included events up to $N+1$ partons in the merging algorithms indicating that it is important to account also for sub-leading jet kinematics when studying jet production in DIS. Comparing the results to experimental data from the H1 collaboration, we find good agreement of differential multi-jet cross sections across a wide range of photon virtualities.
In case of CKKW-L merging we studied two different merging scales, a fixed one and a $Q^{2}$-dependent option referred to as dynamic merging scale. In case of the former the renormalization and factorization scales accounted also for the transverse momentum of the produced partons to partly capture the idea of the dynamic merging scale where the part of the phase space with high-$p_{\mathrm{T}}$ partons and low $Q^{2}$ would be mainly covered by the hard partons instead of parton shower. Indeed, we found that the resulting cross sections were quite similar with these approaches. However, we found that the fixed merging scale lead to an increased sensitivity to the value of the merging scale compared to the dynamic setting and that the resulting $p_{\mathrm{T}}$-slope was slightly steeper. In case of UMEPS we could only use the fixed merging scale option. This algorithm lead to a more stable cross sections against the merging scale variations. However, the agreement with the H1 data was somewhat worse than with CKKW-L for inclusive and dijet production at high $p_{\mathrm{T}}$.
@@ Acknowledgements.
We would like to express our gratitude to Frank Krauss, Leif Lönnblad and Olivier Mattelaer for their insightful discussions and valuable answers regarding this study. We would like to thank Stefan Höche for comments on the DIS merging implementation in Sherpa and help with the scale setting in the HPC-enabled version of Sherpa. We are grateful to Stephen Mrenna for comments on the manuscript. We acknowledge grants of computer capacity from the Finnish Grid and Cloud Infrastructure (persistent identifier urn:nbn:fi:research-infras-2016072533). Part of the computations were carried out on the PLEIADES cluster at the University of Wuppertal, supported by the Deutsche Forschungsgemeinschaft (DFG, grant No. INST 218/78-1 FUGG) and the Bundesministerium für Bildung und Forschung (BMBF).